# Exploring Multiple Strategies to Improve Multilingual Coreference Resolution in CorefUD

### v1

@@ Abstract
Coreference resolution, the task of identifying expressions in text that refer to the same entity, is a critical component in various natural language processing (NLP) applications. This paper presents our end-to-end neural coreference resolution system, utilizing the CorefUD 1.1 dataset, which spans 17 datasets across 12 languages. Our model is based on end-to-end neural coreference resolution system by [1]. We first establish strong baseline models, including monolingual and cross-lingual variations, and then propose several extensions to enhance performance across diverse linguistic contexts. These extensions include cross-lingual training, incorporation of syntactic information, a Span2Head model for optimized headword prediction, and advanced singleton modeling. We also experiment with headword span representation and long-documents modeling through overlapping segments. The proposed extensions, particularly the heads-only approach, singleton modeling, and long document prediction significantly improve performance across most datasets. We also perform zero-shot cross-lingual experiments, highlighting the potential and limitations of cross-lingual transfer in coreference resolution. Our findings contribute to the development of robust and scalable coreference systems for multilingual coreference resolution. Finally, we evaluate our model on CorefUD 1.1 test set and surpass the best model from CRAC 2023 shared task of a comparable size by a large margin. Our nodel is available on GitHub: https://github.com/ondfa/coref-multiling
@@ 1 Introduction
This paper describes our approach to multilingual coreference resolution. The task is based on the CorefUD dataset [3]. The CorefUD 1.1 corpus comprises 17 different datasets spanning 12 languages, all within a harmonized annotation scheme. As CorefUD is intended to extend Universal Dependencies with coreference annotations, all datasets within CorefUD are treebanks. For some languages, dependency annotations were provided by human annotators, while for others, these annotations were generated automatically using a parser. Coreference annotations are built upon these dependencies, meaning that mentions are represented as subtrees in the dependency tree and can be captured by their heads. In some datasets, there are non-treelet mentions—mentions that do not form a single subtree—but even for these, a single headword is selected. Notable differences exist between the datasets, with one of the most prominent being the presence of singletons. Singletons are clusters that contain only one mention and do not participate in any coreference relation, yet they are still annotated as mentions. For further details on the dataset, refer to [3] or [5]. The task was simplified to involve only the prediction of non-singleton mentions and their grouping into entity clusters. Building upon the baseline model proposed by [1], we introduce several novel extensions aimed at enhancing the performance of coreference resolution across multiple languages and datasets within the CorefUD collection. Our primary goal is to develop a universal model capable of handling the diverse and complex nature of these datasets effectively.
@@
Next, we incorporate syntactic information into our model to leverage the dependency structures inherent in the CorefUD datasets. This extension adds depth to the token representations by encoding their paths to the ROOT in the dependency tree, thereby enriching the model’s understanding of syntactic relationships, which are crucial for accurate mention detection and coreference resolution.
In response to the evaluation metrics used in the CRAC 2022 Shared Task, we also propose the Span2head model, which shifts the focus from span-based mention representation to headword prediction. This adjustment aligns the model’s output more closely with the evaluation criteria, optimizing performance by accurately predicting the headwords that serve as minimal spans in coreference chains.
@@
We further address the issue of singletons—mentions that do not participate in coreference chains—by introducing mechanisms to incorporate these into the training process. By modeling singletons explicitly, we ensure that valuable training data from singleton-rich datasets is not discarded, thus improving the model’s robustness and accuracy across different datasets.
Finally, to overcome the limitations of short sequence lengths in models like XLM-R, we propose an innovative approach that utilizes overlapping segments with a cluster merging algorithm. This method ensures that coreference chains spanning multiple segments are correctly identified and merged, even when document segmentation is necessary due to memory constraints.
@@ 2.1 End-to-end Neural Coreference Resolution Models
The first end-to-end neural coreference resolution system was introduced by [6] and many subsequent neural coreference resolution systems are based on their model.
In the model, we start by modeling the probability $P(y_{i}|D)$ of a mention $i$ corefering with the antecedent $y_{i}$ in a document $D$. Since the model adopts the end-to-end approach, the mentions are identified together with the coreference links. We consider every continuous sequence of words as a mention $i$. Therefore, we work with $N=\frac{T(T+1)}{2}$ possible mentions, where $T$ is the number of words in a document $D$.
@@
where $\epsilon$ is an empty (dummy) antecedent. Both scores $s_{m}(i)$ and $s_{a}(i,y_{i})$ are estimated with a feed-forward neural network over the BERT-based encoder.
@@
This end-to-end model has proven to be highly successful, especially on datasets where singletons are not annotated (e.g., CoNLL 2012/OntoNotes), a scenario where traditional two-stage models tend to struggle. The model was later extended by [7] to the higher-order coreference resolution (CR) through iterative refining of span representations with a gated-attention mechanism. They further optimize pro model speed by pruning the mentions with simple scorer at the first step and rescoring only top $k$ spans with the precise scorer and the second step. [8] uses the model from [7] but with BERT as an encoder.
@@ 2.2 Word-level Coreference Resolution
[10] proposed a word-level coreference resolution approach based on the model by [6]. Their method reduces the computational complexity by reducing the mention space. Instead of iterating over all possible spans they first map each gold mention to a single word—specifically, the headword, which they select based on the syntax tree. Antecedent prediction is then performed on the word level. In the next step, they use a span extraction model to predict the original spans from headwords. This model is trained to classify each word in the sentence together with a headword to decide, whether the word is the start token or end token of the span corresponding to the headword. During training the span extraction model is trained simultaneously with with the antecedent prediction as another classification head.
Later [11] proposed a modification in headword selection. They suggested making a coordinate conjunction the head when it appears within an entity to resolve span ambiguity. For example, in the sentence "Tom and Mary are playing," the entities "Tom" and "Tom, and Mary" would typically share the same head in the original approach. Their modification eliminates this ambiguity.
@@ 2.3 CRAC Shared Task on Multilingual Coreference Resolution
CRAC Shared Task on Multilingual Coreference Resolution (CRAC-coref) [12] is an annual shared task that began in 2022 and is built upon CorefUD collection. In fact, this paper is an extension of CRAC22-coref [13] participant system [14], which is based on [15]
Best-performing system in CRAC 2022 was submitted by [16]. It is a two-stage model with joined BERT-like encoder. In mention detection stage, they use extended BIO tagging via stack-manipulation instructions to solve overlapping mentions. Coreference resolution is then done by antecedent prediction head which is very similar to the one use in end-to-end neural CR [6]. They use a concatenation of start and end token as mention representation and classify mention pairs to find the best antecedent of each mention. The schema of CorPipe model is shown in Figure 2.
@@ 3.1.1 Empty Nodes
Some datasets contain empty nodes (also called zeros). These are virtual nodes in the syntactic tree which refers to unexpressed entities. The entities that are not explicitly mentioned in the text, it is implicitly referenced. Empty nodes can make it hard for a general-purpose pre-trained model to understand the language. We cannot discard them from the datasets because most of them are coreferential.
@@ 3.1.2 Singletons
Another key difference between the datasets lies in the annotation of singletons. Singletons are entities with only a single mention throughout the document. In some datasets, singletons are annotated, while in others they are not. That means that when there is an entity with only a single mention, it is not considered an entity during annotation. Table 2 provides the percentage of singleton entities across the datasets (entities with a length of 1).
@@ 3.2 Dataset Lengths
Tables 1, 2, and 3 show basic statistics of all the datasets in CorefUD collection. All these tables are taken from the CorefUD paper [5]. Table 1 shows the basic sizes of the datasets and portion of the train/dev/test split. Column empty shows the number of empty nodes. Table 2 shows statistics about entities. The length of an entity means the number of its mentions thus entities with length 1 are singletons. Table 3 shows statistics of mentions. Here the length of a mention means the numbers of its words. The last three columns shows the percentage of specific mentions. w/empty is mentions with at least one empty node, w/gap is mentions with at least one gap e.g. discontinuous mentions. non-tree are mentions which do not form a single subtree of the syntax tree.
Table 1: Dataset Statistics (taken from [3])
@@
Table 2: Entity Statistics (taken from [3])
@@
Table 3: Mention Statistics (taken from [3])
@@ 3.3 Evaluation Metrics
For evaluation, we employ the official metric for CRAC 2023 Multilingual Coreference Resolution shared task implemented via the CoreUD scorer [^1] https://github.com/ufal/corefud-scorer. The metric is a modification of the standard F1 metric for coreference resolution, averaging three metrics: CLEAF $b^{3}$ and MUC. The main modification of the CorefUD scorer lies in using the head matching. In head matching the gold and system mentions are considered identical if and only if they have the same syntactic head. Another important aspect is that the primary metric ignores singletons. The CorefUD scorer also provides additional metrics, such as BLANC [20] and LEA [21]. To assess the quality of mention matching while ignoring the assignment of mentions to coreferential entities, it uses the MOR score (mention overlap ratio) [13].
@@ 3.3.1 A Link-Based Metric: The muc Score
where $\mathcal{P}(K_{i},R)$ is the partition function, which returns all the partitions of key entity $K_{i}$ with respect to a system’s response $R$. 1 Precision is computed by summing up the correct links in each coreference chain in the response and dividing that by the total number of links in the response—that is, by swapping key and response in the formula above.
@@ 3.3.3 An Entity-Based Metric: ceaf
$B^{3}$ also suffers from a problem—namely, that a single chain in the key or response can be credited several times. This leads to anomalies; for instance, if all coreference chains in the key are merged into one in the response, the $b^{3}$ recall is one. The ceaf metric was proposed by [24] to correct this problem. The key idea of ceaf is to align chains (entities) in the key and response using a map $g$ in such a way that each chain $K_{i}$ in the key is aligned with only one chain $g(K_{i})$ in the response and to then use the similarity $\phi(K_{i},g(K_{i}))$ to compute recall and precision. Because different maps are possible, the one that achieves optimal similarity is used. [25]
@@ 4 Baseline Model & Extensions
We adopted a model from [1] and we use it as a baseline model (without higher-order inference). We use UDAPI [^2] https://github.com/udapi/udapi-python to load the data in CorefUD format. We use XLM-Roberta-large as an encoder.
@@ 4.1 Cross-lingual Training
As you can see from Table 4, there are approximately 50 million parameters trained from scratch for XLM-R. For smaller datasets, it is practically impossible to train so many random parameters. To address this issue, we first pre-train the model on the joined dataset and then fine-tune the model for a specific language.
@@ 4.2 Syntactic Information
We believe that incorporating dependency information can significantly enhance the model, particularly when manually annotated dependencies are available, such as in the Czech PDT dataset. Moreover, dependency information is essential for identifying mention heads, as some datasets in the CorefUD collection annotate mentions as nodes in the dependency tree rather than spans in the linear text flow.
To encode syntactic information, we add to each token representation its path to ROOT in the dependency tree. In more detail, we first set the maximum tree depth parameter and then concatenate Bert representations of all parents up to max depth with the embedding of the corresponding dependency relation. Thus the resulting tree structure representation has the size of $max\_tree\_depth\times(bert\_emb\_size+deprel\_emb\_size)$. This representation is then concatenated with BERT embedding of each token.
@@ 4.3 Span2head Model
Given that the official scorer uses min-span evaluation with headwords as the min spans, we decided to train the model to predict heads rather than entire spans to optimize the evaluation metric. Having all the useful information (even dependency trees), the model should learn the original rules for selecting the head.
@@ 4.5 Singletons
Some datasets in CorefUD collection have singletons annotated and others do not. Specifically, in CorefUD 1.1, 8 out of 17 datasets have more than 10% singletons and 6 of these have more than 70% singletons which is probably a sign of consistent entity annotation independent of the coreference annotation. The baseline model completely ignores singletons during training [^3] The loss is the sum over all correct antecedents and since singletons have no gold antecedents they do not affect the loss. As a result for these 6 singleton-including datasets, we discard more than 70% of training data for mention identification task. To leverage this data, we incorporated singleton modeling into our model and propose several approaches to address this.
@@ 4.5.1 Another Dummy Antecedent
In the first proposed approach, we introduce another virtual antecedent (learned) representing that the mention has no real antecedent, but it is a valid mention, so it is a singleton. For these mentions, we avoid using the binary score because it does not make sense to do it. It would model the similarity between all the singletons (because all the singleton representations would be trained to be similar to the dummy antecedent). Additionally, included an option to use a separate feed-forward network for predicting singleton scores (different from the one for other mentions scores)

### v2

@@ Abstract
Coreference resolution, the task of identifying expressions in text that refer to the same entity, is a critical component in various natural language processing (NLP) applications. This paper presents our end-to-end neural coreference resolution system, utilizing the CorefUD 1.1 dataset, which spans 17 datasets across 12 languages. Our model is based on the end-to-end neural coreference resolution system by [1]. We first establish strong baseline models, including monolingual and cross-lingual variations, and then propose several extensions to enhance performance across diverse linguistic contexts. These extensions include cross-lingual training, incorporation of syntactic information, a Span2Head model for optimized headword prediction, and advanced singleton modeling. We also experiment with headword span representation and long-document modeling through overlapping segments. The proposed extensions, particularly the heads-only approach, singleton modeling, and long document prediction, significantly improve performance across most datasets. We also perform zero-shot cross-lingual experiments, highlighting the potential and limitations of cross-lingual transfer in coreference resolution. Our findings contribute to the development of robust and scalable coreference systems for multilingual coreference resolution. Finally, we evaluate our model on the CorefUD 1.1 test set and surpass the best model from the CRAC 2023 shared task of comparable size by a large margin. Our model is available on GitHub: https://github.com/ondfa/coref-multiling
@@ 1 Introduction
This paper describes our approach to multilingual coreference resolution. The task is based on the CorefUD dataset [3]. The CorefUD 1.1 corpus comprises 17 different datasets spanning 12 languages, all within a harmonized annotation scheme. As CorefUD is intended to extend Universal Dependencies with coreference annotations, all datasets within CorefUD are treebanks. For some languages, dependency annotations were provided by human annotators, while for others, these annotations were generated automatically using a parser. Coreference annotations are built upon these dependencies, meaning that mentions are represented as subtrees in the dependency tree and can be captured by their heads. In some datasets, there are non-treelet mentions—mentions that do not form a single subtree—but even for these, a single headword is selected. Notable differences exist between the datasets, with one of the most prominent being the presence of singletons. Singletons are clusters that contain only one mention and do not participate in any coreference relation, yet they are still annotated as mentions. For further details on the dataset, refer to [3] or [5]. We use the primary evaluation metric from CRAC 2023 Multilingual Coreference Resolution Shared Task. The metric ignores singletons. Building upon the baseline model proposed by [1], we introduce several novel extensions aimed at enhancing the performance of coreference resolution across multiple languages and datasets within the CorefUD collection. Our primary goal is to develop a universal model capable of handling the diverse and complex nature of these datasets effectively.
@@
Next, we incorporate syntactic information into our model to leverage the dependency structures inherent in the CorefUD datasets. This extension adds syntax to the token representations by encoding their paths to the ROOT in the dependency tree, thereby enriching the model’s understanding of syntactic relationships, which are crucial for accurate mention detection and coreference resolution.
In response to the evaluation metrics used in the CRAC 2022 Shared Task, we also propose the Span2head model, which simplifies span-based mention representation to headword prediction. This adjustment aligns the model’s output more closely with the evaluation criteria, where we consider two mentions the same if they share the same head.
@@
We further address the issue of singleton—mentions by introducing mechanisms to incorporate these into the training process. By modeling singletons explicitly, we ensure that valuable training data from singleton-rich datasets is not discarded, thus improving the model’s robustness and accuracy across different datasets.
Finally, to overcome the limitations of short sequence lengths in XLM-R, we propose an approach that utilizes overlapping segments with a cluster merging algorithm. This method ensures that coreference chains spanning multiple segments are correctly identified and merged, even when document segmentation is necessary due to memory constraints.
@@ 2.1 End-to-end Neural Coreference Resolution Models
The first end-to-end neural coreference resolution system was introduced by [6], and many subsequent neural coreference resolution systems are based on their model.
The model estimates the probability $P(y_{i}|D)$ of a mention $i$ corefering with the antecedent $y_{i}$ in a document $D$. Since the model adopts the end-to-end approach, the mentions are identified together with the coreference links. We consider every continuous sequence of words as a mention $i$. Therefore, we work with $N=\frac{T(T+1)}{2}$ possible mentions, where $T$ is the number of words in a document $D$.
@@
where $\epsilon$ is an empty (also referred as dummy) antecedent. If a mention candidate is aligned to the dummy antecedent and it is not an antecedent of any other mention (so the mention should be in the cluster with dummy antecedent only), it is not considered a valid mention. As a result, every valid mention has to be in a cluster with at least one other mention, so the model cannot predict singletons. Both scores $s_{m}(i)$ and $s_{a}(i,y_{i})$ are estimated with a feed-forward neural network over the BERT-based encoder.
@@
This end-to-end model has proven to be highly successful, especially on datasets where singletons are not annotated (e.g., CoNLL 2012/OntoNotes), a scenario where traditional two-stage models tend to struggle. The model was later extended by [7] to the higher-order coreference resolution (CR) through iterative refining of span representations with a gated attention mechanism. They further optimize pro model speed by pruning the mentions with simple scorer at the first step and rescoring only top $k$ spans with the precise scorer and the second step. [8] uses the model from [7] but with BERT as an encoder.
@@ 2.2 Word-level Coreference Resolution
[10] proposed a word-level coreference resolution approach based on the model by [6]. Their method reduces the computational complexity by reducing the mention space. Instead of iterating over all possible spans they first map each gold mention to a single word—specifically, the headword, which they select based on the syntax tree (as the node closest to the root of the tree). Antecedent prediction is then performed on the word level. In the next step, they use a span extraction model to predict the original spans from headwords. This model is trained to classify each word in the sentence together with a headword to decide, whether the word is the start token or end token of the span corresponding to the headword. During training the span extraction model is trained simultaneously with with the antecedent prediction as another classification head.
Later [11] proposed a modification in headword selection, suggesting that when a coordinate conjunction appears within an entity, it should be made the head to resolve span ambiguity. For example, in the sentence "Tom and Mary are playing," the entities "Tom" and "Tom, and Mary" would typically share the same head in the original approach. Their modification eliminates this ambiguity.
@@ 2.3 CRAC Shared Task on Multilingual Coreference Resolution
CRAC Shared Task on Multilingual Coreference Resolution (CRAC-coref) [12] is an annual shared task that began in 2022 and is built upon the CorefUD collection. In fact, this paper is an extension of CRAC22-coref [13] participant system [14], which is based on [15]
Best-performing system in CRAC 2022 was submitted by [16]. It is a two-stage model with a joined BERT-like encoder. In the mention detection stage, they use extended BIO [^1] BIO stands for Beginning, Inside, and Outside. It is a tagging format to label each token of a text as being the beginning (B), inside (I), or outside (O) of an entity. tagging via stack manipulation instructions to solve overlapping mentions. Coreference resolution is then done by antecedent prediction head, a component responsible for identifying the most likely antecedent of each mention, which is very similar to the one used in end-to-end neural CR [6]. They use a concatenation of start and end tokens as mention representation and classify mention pairs to find the best antecedent of each mention. The schema of the CorPipe model is shown in Figure 2.
@@ 3.1.1 Empty Nodes
Some datasets contain empty nodes. These are virtual nodes added to a syntactic tree during the annotation process (and also into the text). These nodes refer to unexpressed entities; entities that are not explicitly mentioned in the text but are referenced. Adding empty nodes can make it hard for a general-purpose pre-trained model to to accurately interpret the meaning of the text. We cannot discard them from the datasets because most of them are part of a coreference cluster. Entity mentions that span only the empty nodes are called zero mentions.
@@ 3.1.2 Singletons
Another key difference between the datasets lies in the annotation of singletons. Singletons are entities with only a single mention throughout the document so these mentions have no coreference relations. In some datasets, singletons are annotated, while in others they are not. That means that when there is an entity with only a single mention, it is not considered an entity during annotation. Table 2 provides the percentage of singleton entities across the datasets (entities with a length of 1).
@@ 3.2 Dataset Lengths
Tables 1, 2, and 3 show basic statistics of all the datasets in CorefUD collection. All these tables are taken from the CorefUD paper [5]. Table 1 shows the basic sizes of the datasets and portion of the train/dev/test split. Column empty shows the number of empty nodes. Table 2 shows statistics about entities. The length of an entity means the number of its mentions thus entities with length 1 are singletons. Table 3 shows statistics of mentions. Here the length of a mention means the number of its words. The last three columns show the percentage of specific mentions. w/empty are mentions with at least one empty node, w/gap are mentions with at least one gap e.g. discontinuous mentions. non-tree are mentions which do not form a single subtree of the syntax tree.
Table 1: Dataset Statistics (taken from [3]). words is the number of non- empty UD nodes. empty is the number of empty UD nodes.
@@
Table 2: Entity Statistics (taken from [3]). The total number of entities and the average number of entities per 1000 tokens in the running text. The maximum and average entity “length”, i.e., number of mentions of the entity. Distribution of entity lengths.
@@
Table 3: Non-singleton mention Statistics (taken from [3]). The total number of mentions and the average number of mentions per 1000 words of running text. The maximum and average mention length, i.e., number of non-empty nodes in the mention. Distribution of mention lengths.
@@ 3.3 Evaluation Metrics
For evaluation, we employ the official metric for CRAC 2023 Multilingual Coreference Resolution shared task implemented via the CoreUD scorer [^2] https://github.com/ufal/corefud-scorer. The metric is a modification of the CoNLL score for coreference resolution, averaging three metrics: CEAF-e, $B^{3}$, and MUC. All these three metrics are described later in this section. The main modification of the CorefUD scorer lies in using the head matching. In head matching the gold and system mentions are considered identical if and only if they have the same syntactic head. Another important aspect is that the primary metric ignores singletons. The CorefUD scorer also provides additional metrics, such as BLANC [20] and LEA [21]. To assess the quality of mention matching while ignoring the assignment of mentions to coreferential entities, it uses the MOR score (mention overlap ratio) [13].
@@ 3.3.1 A Link-Based Metric: The MUC Score
where $\mathcal{P}(K_{i},R)$ is the partition function, which returns all the partitions of key entity $K_{i}$ with respect to a system’s response $R$. Coreference chains are formed by linking each mention in a cluster to its antecedent, creating a sequence of references. Precision is computed by summing up the correct links in each coreference chain in the response and dividing that by the total number of links in the response. Essentially, precision evaluates how many of the predicted coreference links accurately match the gold standard. This is done by swapping key and response in the formula, where the system’s response is compared against the gold coreference chain, and the number of correct links is divided by the total predicted links.
@@ 3.3.3 An Entity-Based Metric: CEAF
$B^{3}$ also suffers from a problem—namely, that a single chain in the key or response can be credited several times. This leads to anomalies; for instance, if all coreference chains in the key are merged into one in the response, the $b^{3}$ recall is one. The ceaf metric was proposed by [24] to correct this problem. The key idea of CEAF is to align chains (entities) in the key and response using a map $g$ in such a way that each chain $K_{i}$ in the key is aligned with only one chain $g(K_{i})$ in the response and to then use the similarity $\phi(K_{i},g(K_{i}))$ to compute recall and precision. Because different maps are possible, the one that achieves optimal similarity is used. [24] present two entity-pair similarity measures that can be used in CEAF: one is the absolute number of common mentions between two entities, and the other is a “local” mention F-measure between two entities. The two measures lead to the mention-based CEAF-m and entity-based CEAF-e, respectively. [25]
@@ 4 Baseline Model & Extensions
We adopt a model from [1] and we use it as the baseline model (without higher-order inference). We use UDAPI [^3] https://github.com/udapi/udapi-python to load the data in CorefUD format. We use XLM-Roberta-large as an encoder.
@@ 4.1 Cross-lingual Training
As you can see from Table 4, there are approximately 50 million parameters trained from scratch for XLM-R. For smaller datasets, it is practically impossible to train so many random parameters from scratch. To address this issue, we first pre-train the model on the joined dataset and then fine-tune the model for a specific language.
@@ 4.2 Syntactic Information
We believe that incorporating syntactic information can significantly enhance the model, particularly when manually annotated dependencies are available, such as in the Czech PDT dataset. Moreover, dependency information is essential for identifying mention heads, as some datasets in the CorefUD collection annotate mentions as nodes in the dependency tree rather than spans in the linear text flow.
To encode syntactic information, we add to each token representation its path to the ROOT in the dependency tree. In more detail, we first set the maximum tree depth parameter and then concatenate Bert representations of all parents up to max depth with the embedding of the corresponding dependency relation. Thus the resulting tree structure representation has the size of $max\_tree\_depth\times(bert\_emb\_size+deprel\_emb\_size)$. This representation is then concatenated with BERT embedding of each token.
@@ 4.3 Span2head Model
Given that the official scorer uses head-match evaluation, we decided to train the model to predict heads rather than entire spans to optimize the evaluation metric. Having all the useful information (even dependency trees), the model should learn the original rules for selecting the head.
@@ 4.5 Singletons
Some datasets in the CorefUD collection have singletons annotated and others do not. Specifically, in CorefUD 1.1, 8 out of 17 datasets have more than 10% singletons and 6 of these have more than 70% singletons which is probably a sign of consistent entity annotation independent of the coreference annotation. The baseline model completely ignores singletons during training [^4] The loss is the sum over all correct antecedents and since singletons have no gold antecedents they do not affect the loss.. As a result for these 6 singleton-including datasets, we discard more than 70% of training data for mention identification task. To leverage this data, we incorporate singleton modeling into our model and propose several approaches to address this.
@@ 4.5.1 Another Dummy Antecedent
In the first proposed approach, we introduce another virtual antecedent (with trained embedding) representing that the mention has no real antecedent, but it is a valid mention, so it is a singleton. For these mentions, we avoid using the binary score because it would unnecessarily model the similarity between all the singletons 555because all the singleton representations would be trained to be similar to the dummy antecedent. Additionally, included an option to use a separate feed-forward network for predicting singleton scores (different from the one for other mentions scores).